OpenStudy (anonymous):

Frances has 4 square baking dishes. Each pan is 7 inches × 7 inches × 2 inches. Which expression will give her the total volume of all the pans?

OpenStudy (anonymous):

A. 7^2 × 2 B. 7 × 2^4 C. (7 × 2 × 2) × 4 D. (7^2 × 2) × 4

OpenStudy (unklerhaukus):

volume per dish is length x width x height = 7x7x2 total Volume = 4 dishes x volume/dish so \[4 \times 7 \times 7 \times 2\] which is equivalent to answer D.